Circular No. 9103

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2009 Y2 (KOWALSKI) As first announced on CBET 2088, R. A. Kowalski (Lunar and Plan-

etary Laboratory) reports a cometary object (discovery observation tabulated below) found on images taken with the Catalina Sky Survey’s 0.68-m Schmidt telescope; in good seeing, it appeared more diffuse than a typical object of comparable brightness and showed what appeared to be a round coma with strong condensation in p.a. 270^{o}. Observations by A. R. Gibbs with the Mt. Lemmon 1.5-m reflector on Dec. 20.2 UT showed a 4′′.5 broad tail in p.a. 45^{o}. After the discovery had been added to the Minor Planet Center’s ‘NEO Confirmation Page’, W. H. Ryan and E. V. Ryan (Magdalena Ridge Observatory, 2.4-m f/8.9 reflector) on Dec. 20.2 and 21.1 confirmed a faint tail in p.a. 45^{o}. A 105-min integration by P. Birtwhistle (Great Shefford, Berkshire, England, 0.40-m f/6 Schmidt-Cassegrain reflector) on Dec. 20.9 showed the object to have a diameter of ∼ 6′′ with a 10′′ extension in p.a. 50^{o}; a faint, straight tail was suspected, also in p.a. 50^{o}, at least 25′′ (and possibly up to 45′′) long.

2009 UT α_{2000} δ_{2000} Mag. Observer
Dec. 20.11774 00^{h}37^{m}55.^{s}48 +31^{o}27′27′′.2 18.9 Kowalski

The initial astrometry, along with preliminary parabolic orbital elements, appeared on MPEC 2009-Y40. Additional astrometry has led to the computation of the following elliptical elements (from MPC 67973):

T = 2010 Mar. 30.696 TT  ω = 171.^{o}805
e = 0.62349  Ω = 262.256 }2000.0
q = 2.33726 AU  i = 29.864

a = 6.20772 AU n^{o} = 0.063724 P = 15.47 years

η CARINAE

A. Verveer, Perth Observatory; and D. Frew, Macquarie University, report the following B and V photometry of η Car obtained with a 40-cm f/11 reflector (+ Apogee AP7 CCD camera, using standard Bessell filters; comparison star HDE 303308) at Perth Observatory: June 5.526 UT, V = 4.72, B–V = +0.71; Nov. 20.7914, 4.665, +0.709; 21.7733, 4.655, +0.714; 25.7814, 4.665, 0.714. The photometry was summed through a 22′′ aperture and naturally includes the central binary star plus all of the surrounding Homunculus nebula; the recent HST photometry of Davidson et al. (IAUC 9094) refers only to the star itself and is about a half magnitude fainter.

2009 December 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green